A gas sample used to simulate the atmosphere of an exoplanet contains 7.50 x 102 mg...

  1. A gas sample used to simulate the atmosphere of an exoplanet contains 7.50 x 102 mg N2 and 1.25 x 102 mg CH4. The partial pressure of CH4 at 300.0 K is 4.5 kPa. What volume does the entire gas sample occupy?
    1. 68.9 L
    2. 0.550 L
    3. 4.32 L
    4. 34.0 L
    5. 0.693 L
Answer #1

no of moles of N2   = W/G.M.Wt

                              = 0.75/28   = 0.0268moles

no of moles of CH4   = W/G.M.Wt

                                 = 0.125/16   = 0.0078125moles

total no of moles of N2 and CH4 = nN2 + nCH4

                                                    = 0.0268+0.0078125

                                                    = 0.0346moles

mole fraction of CH4   = nCH4/nCH4 + nN2

                                  = 0.0078125/0.0346 = 0.226

partial pressure of CH4   = mole fraction of CH4* total pressure

   4.5                                  = 0.226*total pressure

total pressure   = 4.5/0.226   = 19.9Kpa

P   = 19.9Kpa

      = 19.9*0.00986923                [ 1KPa   = 0.00986923atm]

       =0.1964atm

T     = 300K

n = 0.0346moles

PV   = nRT

V   = nRT/P

      = 0.0346*0.0821*300/0.1964

       = 4.32L

c. 4.32L >>>answer